A particle of mass 1.0 g is moving in a one-dimensional box of length 1.0 cm with a velocity 1.0 cm/s. Find its quantum number n.
Answer #1

THE KINETIC ENERGY OF PARTICLE IS EQUAL TO THE TOTAL ENERGY

m=.001kg,v=.01m/s,l=.01m

kinetic energy =1/2mv2

=1/2*.001*.01*.01

=5*10-8J

NOW THIS ENERGY MUST BE EQUAL TO THE ENERGY OF A PARTICLE IN ONE D BOX WHICH IS GIVEN BY

E=n2h2/8ml2

substituting values

5*10-8=n2h2/8ml2

n2 =5*10-8*8*.001*.01*.01/(6.26*10-34)2

FINALLY THE VALUE OF n is comes out to be 3*1026